Define the relation of d on A by xdy if x is contained within y. For
example, 01d101. Draw a digraph for this relation.
1
Expert's answer
2020-10-29T15:34:42-0400

 Let a, "b \in Z" . a b if and only if there exists an integer k such that "a k = b."



The notation "(4, 16) \in r" or "(3, 7.2) \in s" makes sense in both cases. However, r would be more naturally expressed as "r(x) = x^2" or "r(x)= y" , where "y = x^2" . But this notation when used for s is at best awkward. The notation x § y is clear and self-explanatory; it is a better notation to use than "(x,y)\in s" . Many of the relations we will work with "resemble" the relation "\leq" , so "x s y" is a common way to express the fact that x is related to y through the relation s.